OpenStudy (anonymous):

A mortar shell is fired with horizontal velocity of 300 m/s form the top of cliff 100 m high. (i) How long will it take to reach the ground ? (ii) How far from the foot of cliff will it strike the ground ?

OpenStudy (king):

time=60 s?

OpenStudy (king):

18000 m?

OpenStudy (king):

is it rite?

OpenStudy (king):

hasnain?

OpenStudy (anonymous):

ooh sorry , its is wrong , i have done mistake sorry , i am writting it now

OpenStudy (anonymous):

A mortar shell is fired at a ground level target 490m away with initial velocity of 98m/s. Find the two possible values of the launch angle. Calculate the minimum time to hit the target.

OpenStudy (amistre64):

assuming an unreal parabolic path .... we need to calculate this equation; which turns into the physics equation for it in the end that I can never remember :) \[-\frac{1}{2}g\ t^2+V_i sin(a)\ t + H_i=0\] \[-\frac{1}{2}9.8\ t^2+98 sin(a)\ t + 0=0\] we should know that the outward speed is: 98 cos(a); and that it needs to cover a distance of 480 in t sec. 98 cos(a) t = 480 t = \(\cfrac{480}{98cos(a)}\) lets use this "value" for t in out equation: \[-\frac{1}{2}9.8\ \left(\cfrac{480}{98cos(a)}\right)^2+98 sin(\theta)\ \cfrac{480}{98cos(a)}=0\] \[-\frac{480^2(4.9)}{98^2}sec^2(a)+ 480\ tan(a)=0\] \[-\frac{480^2(4.9)}{98^2}(tan^2(a)+1)+ 480\ tan(a)=0\] \[-\frac{480^2(4.9)}{98^2}tan^2(a)+ 480\ tan(a)-\frac{480^2(4.9)}{98^2}=0\] this of course is a quadratic equation that can be solved for tan(a)

OpenStudy (amistre64):

\[tan(a)=\frac{-480\pm \sqrt{480^2-\cfrac{4(480^4)(4.9^2)}{98^4}}}{\cfrac{-2(480^2)(4.9)}{98^2}}\] that looks fun lol ... better let the wolf handle it :) tan(a) = .26166 and 3.82167 this means that: a = arctan(.26166) or arctan(3.82167) a = 14.66 or 75.34 degrees but after all that youd have to dbl chk it

OpenStudy (amistre64):

and of course I used 480 as my standard typo to mess it all up; ... gotta get to class anyhoos so good luck :)